OpenStudy (anonymous):

A hard rubber rod with an electric potential energy of 5.2 × 10–3 J has a charge of 4.0 µC at the tip. What is the electric potential at the tip?

OpenStudy (irishboy123):

\[V = \frac{U}{Q}\]potential is potential energy per coulumb. the charge is giving rise to the potential and the potential energy.

OpenStudy (michele_laino):

we can think to your system as a capacitor, like this:  now, if I call with U the electrical potential energy, then we can write: \[\Large U = \frac{1}{2}C{V^2} = \frac{1}{2}QV\] where C is the capacitance of your capacitor, and V is the requested voltage. So we have: \[\Large V = \frac{{2U}}{Q} = ...?\]